OpenStudy (anonymous):

the minimum value of sec^2A + cosec^2A is a)1 b) -2 c)2 d)4

OpenStudy (shubhamsrg):

AM>=GM should help your case

OpenStudy (anonymous):

can you explain it more !

OpenStudy (shubhamsrg):

(sec^2 a + cosec^2 a) >= 2 seca cosec this much should be clear ?

OpenStudy (anonymous):

so that means \[4+( \tan \theta - \cot \theta)^2 \ge 4 \]

OpenStudy (anonymous):

and the minimum value be 4 . m i right ?

OpenStudy (shubhamsrg):

how did you get that? :O

OpenStudy (anonymous):

\[1 +\tan^2 \theta + 1+ \cot^2 \theta \]

OpenStudy (anonymous):

\[2 +(\tan \theta -\cot \theta )^2 + 2 \tan \theta .\cot \theta \]

OpenStudy (shubhamsrg):

what after that ?

OpenStudy (anonymous):

then as you have told me \[4+( \tan \theta - \cot \theta )^2\ge 4\]

OpenStudy (shubhamsrg):

how you get 4 on RHS ?

OpenStudy (anonymous):

cause simply you get \[4+ (\tan \theta - \cot \theta )^2 \]

OpenStudy (anonymous):

which is grearter than 4

OpenStudy (shubhamsrg):

ohh.. so you have not used AM>=GM ..oh okay, I wondered how you got that from AM>=GM hmm nice yes, 4 should be the ans

OpenStudy (anonymous):

well thank you !

ganeshie8 (ganeshie8):

nice :)

ganeshie8 (ganeshie8):

another way sec^2A + cosec^2A 1/cos^2A + 1/sin^2A sin^2A+ cos^2A ---------------- cos^2A sin^2A 1 ------------- cos^2A sin^2A 4 ----------- sin^2 2A smallest value = when bottom is large = 4